# Deep-learning-based continuous attacks on QKD protocols

### v3

@@ III.3 Attack via continuous measurement
Figure 7: Mean accuracy $A$ and standard deviation of the neural network as a function of the measurement length $\gamma_{D}\Delta t$. The photo currents were obtained using Eq. (3) and (4). Other parameters are $\gamma_{D}t_{f}=3$ and $\omega=\gamma_{E}=\gamma_{D}$.
III.4 Attack via continuous measurement and quantum feedback
We briefly investigate here the possibility for the spy to apply quantum feedback based on their measurement [51] with the aim of covering their track. We thus adapt Eq. (3) to include the action of a feedback which yields [51]
$d\rho_{J}=$$\big{\{}-i\left[H,\rho_{J}\right]+\gamma_{E}\mathcal{D}[e]\rho_{J}-i\gamma_{E}\left[f,e\rho_{J}+\rho_{J}e^{\dagger}\right]\big{\}}dt$ (16)
$+\gamma_{E}\mathcal{D}[f]\rho_{J}\frac{dt}{\eta}+\gamma_{E}\mathcal{H}\left[\sqrt{\eta}e-if/\sqrt{\eta}\right]\rho_{J}dW$
$+\gamma_{D}\mathcal{D}[d]\rho_{J}dt,$
where $f$ is the feedback operator. The non-selective evolution equation of the system can be obtained by taking the ensemble average of this equation, which removes the term containing $dW$. By doing so we obtain the homodyne mediated feedback master equation in the Lindblad form [74]:
$\dot{\rho}_{J}=$ $-i[H+\frac{\gamma_{E}}{2}(e^{\dagger}f+fe),\rho_{J}]+\gamma_{D}\mathcal{D}[d]\rho_{J}$ (17)
$+\gamma_{E}\mathcal{D}[e-if]\rho_{J}+\frac{1-\eta}{\eta}\gamma_{E}\mathcal{D}[f]\rho_{J}.$
The effect of feedback is to replace the $e$ operator by $e-if$, to add an extra term to the Hamiltonian and a term vanishing for the efficiency $\eta=1$.
To find the optimal feedback operator we parameterize $f$ as
$f=\cos(\phi)\sigma_{x}+\sin(\phi)\sigma_{z}.$ (18)
We also set the measurement and the feedback to be performed between $\gamma_{D}t=0.1$ and $0.5$ (the optimal duration found in the previous section). The QBER as a function of the measurement angle $\theta$ and the feedback angle $\phi$ is displayed in Fig. 8 for $\gamma_{D}t_{f}=3$. For the optimal measurement angle $\theta=1.86\pi$ found earlier, the feedback angle minimizing the QBER of Fig. 8 is $\phi=0.94\pi$. With that angle, we obtain a QBER of $29.9\%$, which is greater than without feedback, which could be due to the fact that an unconditional feedback introduces more noise in the system than it prevents. From this analysis, it seems that quantum feedback does not lead to any improvement of the attack, but many other parameter regimes and feedback schemes, such as feedback conditioned on the measurement record, could be explored in the future.
Dissipation Attack Impact (QBER) Accuracy
@@ III.4 Attack via continuous measurement and quantum feedback
$d=\sigma_{x}$ $e=0$ $25\%$ $0.0$
$d=\sigma_{x}$ Projective measurement $37.5\%$ $69.4\%$
Continuous measurement
$d=\sigma_{x}$ $e=\sigma_{z}$ $49\%$ $72\%$
$d=\sigma_{x}$ $e=0.90\sigma_{x}-0.43\sigma_{z}$ $27.6\%$ $86.1\%$
($\theta=1.86\pi$) ($2.6\%$ increase)
Time Shift $1-2\%$ increase $60-70\%$
Table 2: Summary of the QBER and accuracies $A$ generated by the different attack schemes we analyzed. We compare our results with the Time-Shift attack [17, 75], even though the noise model we use is different. The parameters are $\gamma_{D}t_{f}=3$, $\gamma_{D}\Delta t=0.4$, $\gamma_{D}t^{*}=0.3$ and $\omega=\gamma_{E}=\gamma_{D}$. We set Eve measurement time $\gamma_{D}t^{*}$ to $0.3$ such that it corresponds to the middle of the optimized continuous measurement.
Figure 8: QBER$(\theta,\phi)$ as a function of the measurement and feedback angles: $\theta$ and $\phi$. The evolution of the photon states was obtained using Eq. (17). Other parameters are $\gamma_{D}t_{f}=3$ and $\omega=\gamma_{E}=\gamma_{D}$.
@@ IV.1 Information-Disturbance Principle
[…]
Let $S$ be a discrete variable representing the different initial states ($s=0,1,2,3$) and $X\in\mathbb{R}^{40}$ a $40$-dimensional continuous variable representing the values of the homodyne currents at each time step, we obtain (see Appendix G) that the information gain yielded by the homodyne measurement on the qubits of the BB84 protocol is
$I(S;x)$ $=0.2894\text{ nats}=0.4175\ \text{bits}.$ (22)
At the same time, the estimated bit and phase error rates being, from Eq. (15), $e_{x}=0.275$ and $e_{z}=0.275$, Eq. (19) becomes
@@
If we now consider a regime below the $11\%$ threshold and analyze the information gained by the homodyne measurement when no dissipation is occurring in the optical fiber (such that all of the QBER is caused by the spy), we obtain (with $\gamma_{D}=0$, $\gamma_{E}=\omega/3$ and $2\times 10^{4}$ trajectories)
$\mathrm{QBER}=5.8\%\leq 11\%,$ (24)
$A=88.8\%,$ (25)
$I(S;x)=0.429\ \mathrm{bits}\approx I_{Eve}$ (26)
$I_{Eve}\leq H_{2}(0.058)+H_{2}(0.058)=0.64\ \mathrm{bits}.$ (27)
@@
where $I(A;E)$ is the mutual information between Alice and Eve and $\phi[z]\equiv(1+z)\ln(1+z)+(1-z)\ln(1-z)$. It has been shown that such a bound can be saturated using a phase-covariant quantum cloner, which is an approximate cloning procedure for two-level systems on the equator of the Bloch sphere [53], and that a secure key can be distilled as long as the QBER is below $14.65\%$, point at which the curves $I(A;E)$ and $I(A;B)$ intersect. Inserting this value into Eq. (28), we obtain that Eve’s maximum obtainable information is $0.399$ bits. The analysis above showed slightly better performances for the DLCA (even with a lower QBER): 0.4175 bits with dissipation [Eq. (22)] and 0.429 bits without dissipation [Eq. (26)]. This could be explained by the non-unitary nature of our attack.
@@ IV.2 Key rates
By using the dissipator (32) with $\gamma_{D}=\omega/100$ and taking the optimal angle and measurement duration found above, Eve’s neural network accuracy about Alice’s key becomes $87.8\%$ and the QBER $10.9\%$. Thus, the conditional entropies, per bit, of Alice’s bit given Eve’s and Bob’s are
$H_{2}(A\mid E)$ $=H_{2}(0.878)=0.535\ \mathrm{bits},$ (33)
$H_{2}(A\mid B)$ $=H_{2}(1-0.109)=0.497\ \mathrm{bits},$ (34)
@@
Time-shift attack.
For comparison, the Time-Shift attack mentioned above would give
$R=H_{2}(0.7)-H_{2}(0.067)=0.57\ \mathrm{bits},$ (36)
with an accuracy of $70\%$ and a QBER increase of $1\%$ (best values in Table 2). These results show that while our attack does not question the security of the protocol, it does affect the minimum number of resources (and thus the maximum key rate) Alice and Bob could use to secure the protocol.
@@ V Conclusion
In this paper, we introduced a new type of individual attack on QKD protocols based on continuous measurement that, used as an input of a trained recurrent neural network, allows the spy to retrieve with high accuracy the initial states sent by one of the parties without being significantly noticed. We denote our attack as a Deep-learning-based continuous attack (DLCA). Although more quantitative and comparison analyses should be done, also in terms of noise models considered for the optical fiber, our attack scheme exhibits performances similar to the optimal individual attack.
To go further, one could investigate the possible generalization of our strategy to a collective or coherent attack [78].
Also, a more complex and realistic noise model of the optical fiber could be used. In [26], Kozubov et al. span the space using three states: the vacuum state and the states we denoted |$\ket{0}$ and |$\ket{1}$ in this work (i.e., horizontally and vertically polarized photons). By doing so, they take into account the non-zero probability that the photon is absorbed in the optical fiber. They also tune the phenomenological parameters involved in the master equation, which allows to take into account the phenomena of birefringence, isotropic absorption, and dichroism. One could also consider the potential losses caused by the imperfection of Alice and Bob detectors. Overall, it is straightforward to adapt our approach to such other noise models. Note that our attack does not compromise the security of the BB84 protocol as long as Alice and Bob perform enough privacy amplification, reducing the key rate at least below the upper bound we computed.
@@
Also, we considered here the introduction of unconditional quantum feedback. However, one could consider feedback conditioned on the measurement outcome history and thus implement conditional feedback, yielding Non-Markovian dynamics. Also, one could describe the dynamics of the monitored system (without feedback) with a Non-Markovian approach, using for example the cHEOM method [93].
@@ Appendix B Analytical derivation of the QBER without attacks
where the jump operator $L=\sigma_{x}$ models bit flip errors, and the Hamiltonian is $H=\omega\sigma_{z}$ for the initial states |$\ket{0}$ and |$\ket{1}$ and $H=\omega\sigma_{x}$ for the initial states $\ket{+}$ and $\ket{-}$. We denote the matrix elements of $\rho$ in the basis {|$\{\ket{0},\ket{1}\}$ as $\rho_{ij}=Tr\left(\ket{j}\bra{i}\rho\right)=\bra{i}\rho\ket{j}(i,j=0,1)$. Projecting the master equation in the computational basis gives the following linear set of equations for the density matrix elements
$\begin{cases}\dot{\rho}_{00}=\gamma_{D}(\rho_{11}(t)-\rho_{00}(t))\\ \dot{\rho}_{01}=\gamma_{D}(\rho_{10}(t)-\rho_{01}(t))\\ \dot{\rho}_{10}=\gamma_{D}(\rho_{01}(t)-\rho_{10}(t))\\ \dot{\rho}_{11}=\gamma_{D}(\rho_{00}(t)-\rho_{11}(t))\\ \end{cases}$ (41)
@@
so that the probability that the qubit is in the state |$\ket{0}$ after going through the optical fiber is $\rho_{00}(t)=\frac{e^{-2\gamma_{D}t}}{2}+\frac{1}{2}$.
In the case $\rho(0)=\ket{1}\bra{1}$ |, we find
@@
and the probability that the qubit is in the state |$\ket{1}$ is given by $\rho_{11}(t)=\frac{e^{-2\gamma_{D}t}}{2}+\frac{1}{2}$.
In the case $\rho(0)=\ket{+}\bra{+}$ |, we find
@@
Finally, in the case $\rho(0)=\ket{-}\bra{-}$ |, we find
@@
As the states $\ket{+}$ and $\ket{-}$ are eigenstates of the $\sigma_{x}$ jump operator, they will not change when traveling through the optical fiber: they are insensitive to the dissipation process. These states are decoherence-free states (or dark states) for the given master equation.
@@ Appendix C Analytical derivation of Eve accuracy for a projective measurement
The result above allows us to easily determine Eve accuracy in the case she performs a projective measurement at time $t^{*}$. Indeed, when the photon reaches Eve at time $t^{*}$, she also chooses randomly one of the two available bases. The probabilities above correspond to someone measuring in the right basis. However, if Eve does not, her probability of detecting the right state is still $1/2$ is still 1/2 as she can get each result with equal probability. Thus, the probability that Eve deduces the right state (i.e., the accuracy $A$) can be obtained from Eq. (47), which yields
@@ Appendix D Analytical derivation of Bob accuracy for a projective measurement (intercept-and-resend attack)
After Eve’s projective measurement at time $t^{*}$, the photon is in the state she measured, and thus evolves according to Eq. (42) until it reaches Bob at time $t_{f}$. To simplify the process, we will look in detail to the case where Alice sends the initial state |$\ket{0}$ which generalizes easily to the other states. Since we want to obtain the probability that Bob measures the state Alice sent, which is $1-\mathrm{QBER}$, and since they will both, at some part of the protocol, compare the bases they respectively used and discard the differing ones (see Sec. II.1), we consider that Bob measures in the Pauli-Z basis {|$\{\ket{0},\ket{1}\}$. There are four distinct cases, each corresponding to a different measurement result for Eve.
If she measures in the Pauli-Z basis (probability $1/2$) and she measures the state |$\ket{0}$ [probability $(1+e^{-2\gamma_{D}t^{*}})/2$], the photon will be in the state |$\ket{0}$ right after. Thus Bob will measure the state |$\ket{0}$ with probability $(1+e^{-2\gamma_{D}(t_{f}-t^{*})})/2$.
However, if Eve measures in the Pauli-Z basis but the result is |$\ket{1}$ [probability $(1-e^{-2\gamma_{D}t^{*}})/2$] then the probability that Bob measures the state |$\ket{0}$ is $(1-e^{-2\gamma_{D}(t_{f}-t^{*})})/2$.
If Eve measures in the Pauli-X basis (probability $1/2$), the result will be either $\ket{+}$ or $\ket{-}$, each with probability $1/2$, which is the state that will reach Bob since they are not affected by the dissipation. Therefore, Bob’s result will be |$\ket{0}$ or |$\ket{1}$, each with probability $1/2$.
Altogether, this yields the probability that Bob measures the state |$\ket{0}$ if Alice sent it
$\text{I\kern-1.49994ptP}_{b}(\ket{0})$ $=\frac{1}{2}\left(\frac{1}{2}+\frac{e^{-2\gamma_{D}t^{*}}}{2}\right)\left(\frac{1}{2}+\frac{e^{-2\gamma_{D}(t_{f}-t^{*})}}{2}\right)$ (50)
@@
$\text{I\kern-1.49994ptP}_{b}(\ket{1})$ $=\frac{2}{4}+\frac{e^{-2\gamma_{D}t_{f}}}{4},$ (51)
$\text{I\kern-1.49994ptP}_{b}(\ket{+})$ $=\text{I\kern-1.49994ptP}_{b}(\ket{-})=\frac{1}{2}\times 1\times 1+2\times\frac{1}{2}\times\frac{1}{2}\times\frac{1}{2}$
@@
$\text{I\kern-1.49994ptP}_{b}(\text{right result})$ $=\frac{1}{4}\text{I\kern-1.49994ptP}_{b}(\ket{0})+\frac{1}{4}\text{I\kern-1.49994ptP}_{b}(\ket{1})$ (53)
$+\frac{1}{4}\text{I\kern-1.49994ptP}_{b}(\ket{+})+\frac{1}{4}\text{I\kern-1.49994ptP}_{b}(\ket{-})$
@@
Appendix E Confidence of the model for $\sigma_{z}$ and $\sigma_{x}$ measurements
We compute the confidence of our model by summing the probability that the network outputs for the ground truth state (i.e., the initial state it must predict). The results are displayed in Fig. 9(a) and 9(b) for the measurement operators $e=\sigma_{z}$ and $e=\sigma_{x}$ respectively. For $\sigma_{z}$, the network is correct almost every time for the states |$\ket{0}$ and |$\ket{1}$ while for the two others it either only predicts $\ket{+}$ or $\ket{-}$ depending on the first it has seen during training, which varies since the batches order is random. However, these four trainings yield similar accuracies. This result confirms our intuition that the model cannot distinguish the states $\ket{+}$ and $\ket{-}$ when setting the measurement operator to $e=\sigma_{z}$, thus performing random guess. On the other hand for $\sigma_{x}$, we see that the network much more often predicts one of the four states (the first it has seen during training) and almost never outputs one or several of them.
The architecture is robust to changes in the training order for the Pauli-Z measurement operator while it is not for the Pauli-X one, which is due to the fact that the mean accuracy achieved with the latter is low ($50\%$) while the standard deviation is high ($15\%$), meaning that the network has trouble learning something meaningful from the dataset and is thus unstable.
Figure 9: Comparison of the number of right predictions, ground truth and of the model confidence about them for each initial states, for four different training. These results are obtained from a measurement performed with the operator (a) $e=\sigma_{z}$ and (b) $e=\sigma_{x}$. In panel (a), the network is correct almost every time for the states |$\ket{0}$ and |$\ket{1}$ while for the two others it either only predicts $\ket{+}$ or $\ket{-}$ depending on the first it has seen during training (order of the batches is random). These four trainings yielded accuracies around $71\%$. In panel (b), we see that the network much more often predicts one of the four states (the first it has seen during training) and almost never outputs one or several of them. The four trainings yielded accuracies of $25\%$, $34\%$, $45\%$ and $63\%$ respectively (going from left to right and top to bottom).
Appendix F LSTM hidden states activation through time
For the initial state |$\ket{0}$ (panel (a) of Fig. 10), almost all hidden states of the LSTM network activate at the early times of the time-series before maintaining a non-zero activity during the whole time sequence. However, for the initial state $\ket{+}$ (panel (b) of Fig. 10), the hidden states that activate retain their activity until the end of the time series, which could indicate that more information about the initial state is contained in the later times of the photo currents. The early activation of the hidden states in both cases demonstrates that most information is contained in the first time steps of the currents, and is memorized by the hidden units until the output.
Figure 10: LSTM hidden states activations through the input time steps for the initial state |$\ket{0}$ [panel (a)] and $\ket{+}$ [panel (b)].
Appendix G Information gain computation

### v4

@@ III.3 Attack via continuous measurement
Figure 6: Mean accuracy $A$ and standard deviation of the neural network as a function of the measurement length $\omega\Delta t$. The photo currents were obtained using Eq. (3) and (4). Other parameters are $\omega t_{f}=0.1$ and $\gamma_{E}=5\omega=5\gamma_{D}$.
Dissipation Attack Impact (QBER) Key accuracy
@@
$d=\sigma_{x}$ $e=0$ $4.5\%$ $0.0$
$d=\sigma_{x}$ Projective measurement $27.2\%$ $74.1\%$
DLCA
$d=\sigma_{x}$ $e=\sigma_{z}\ (\gamma_{E}=5\omega)$ $20.5\%$ $68\%$
$d=\sigma_{x}$ $e=-\sigma_{x}\ (\gamma_{E}=10\omega)$ $20\%$ $73\%$
($\theta=\pi$)
Table 2: Summary of the QBER and accuracies $A_{key}$ generated by the different attack schemes we analyzed. The parameters are $\omega t_{f}=0.1$, $\omega\Delta t=0.07$, and $\omega=\gamma_{D}$. We set Eve measurement time $\omega t^{*}$ to $0.35$ such that it corresponds to the middle of the optimized continuous measurement.
@@ IV.1 Information-Disturbance Principle
$I(X;Y)=\iint P_{X,Y}(x,y)\log\left(\frac{P_{X,Y}(x,y)}{P_{X}(x)P_{Y}(y)}\right)dx\,dy.$ (18)
Let $S$ be a discrete variable representing the different initial states ($s=0,1,2,3$) and $X\in\mathbb{R}^{70}$ a $70$-dimensional continuous variable representing the values of the homodyne currents at each time step, we obtain (see Appendix E) that the $95\%$ confidence interval on the information gain yielded by the homodyne measurement, with the parameters of Eq. (14) and (15), on the qubits of the BB84 protocol is
$I(S;x)=\left[0.1519,\ 0.1823\right]\ \text{bits},$ (19)
which represents information about the initial qubit states in $\left[0.3880,\ 0.4275\right]$ bits. Also, one can compute the mutual information between Eve’s and Alice’s keys from the confusion matrix of the deep learning model (see Appendix E), and obtain $I(A;E)=0.1527$ bits. Thus with $95\%$ confidence, the model saturates the data processing inequality, proving the only limitation of the DLCA attack arises from the homodyne measurement itself and the information it extracts.
At the same time, the estimated bit and phase error rates being, from Eq. (15), $e_{x}=0.20$ and $e_{z}=0.20$, Eq. (16) becomes
@@
If we now consider a regime below the $11\%$ threshold and analyze the information gained by the homodyne measurement when no dissipation is occurring in the optical fiber (such that all of the QBER is caused by the spy), we obtain (with $\gamma_{D}=0$ and $\gamma_{E}=4\omega$)
$\mathrm{QBER}=10.7\%\leq 11\%,$ (21)
$A_{key}=69\%,$ (22)
$I(S;x)\in\left[0.0919,\ 0.1222\right]\ \mathrm{bits}\approx I_{Eve}$ (23)
$I_{Eve}\leq H_{2}(0.107)+H_{2}(0.107)=0.98\ \mathrm{bits}.$ (24)
@@
where $I(A;E)$ is the mutual information between Alice and Eve and $\phi[z]\equiv(1+z)\ln(1+z)+(1-z)\ln(1-z)$. It has been shown that such a bound can be saturated using a phase-covariant quantum cloner, which is an approximate cloning procedure for two-level systems on the equator of the Bloch sphere [53], and that a secure key can be distilled as long as the QBER is below $14.65\%$, point at which the curves $I(A;E)$ and $I(A;B)$ intersect. Inserting this value into Eq. (25), we obtain that Eve’s maximum obtainable information is $0.399$ bits. With such QBER, the DLCA attack reaches $A=45.5\%,\ A_{key}=70.7\%$, and the homodyne measurement estimated information gain lies in $\left[0.1204,0.1511\right]$ bits with $95\%$ confidence. Also, the mutual information between Eve’s and Alice’s key is $0.127$ bits.
@@ IV.2 Key rates
By using the dissipator (29) with $\gamma_{D}=\omega/100$, $\gamma_{E}=7\omega/5$ and taking the optimal angle and measurement duration found above, Eve’s neural network accuracy about Alice’s key becomes $68\%$ and the QBER $10.9\%$. Thus, the conditional entropies, per bit, of Alice’s bit given Eve’s and Bob’s are
$H_{2}(A\mid E)$ $=H_{2}(0.68)=0.90\ \mathrm{bits},$ (30)
$H_{2}(A\mid B)$ $=H_{2}(1-0.11)=0.497\ \mathrm{bits},$ (31)
@@ V Conclusion
In this paper, we introduced a new type of individual attack on QKD protocols based on continuous measurement that, used as an input of a trained recurrent neural network, allows the spy to retrieve with high accuracy the sifted key bits sent by one of the parties without being significantly noticed. We denote our attack as a Deep-learning-based continuous attack (DLCA). Although more quantitative and comparison analyses should be done, also in terms of noise models considered for the optical fiber, our attack scheme exhibits better performances than a projective measurement attack. Note that since we assume the use of perfect single photon sources and detectors, our attack does not compromise the security of the BB84 protocol as long as Alice and Bob perform enough privacy amplification, reducing the key rate at least below the upper bound we computed. However, our attack could be adapted to quantum hacking setups targeting vulnerabilities in the implementation of QKD protocols likely to compromise their security. Our work constitutes a first step towards this goal, as well as other promising research directions outlined below.
Indeed, to go further, one could for example investigate the possible generalization of our strategy to a collective or coherent attack [76].
[…]
@@
Also, one could consider that Eve uses quantum feedback based on the measurement outcomes to try to cover her tracks. Depending on the noise model and the regime of parameters, such a conditional feedback could yield Non-Markovian dynamics which could be studied via a Non-Markovian approach such as cHEOM [91].
@@ Appendix B Analytical derivation of the QBER without attacks
[…]
@@
[…]
In the case $|\rho(0)=\ket{1}\bra{1}$, we find
@@
[…]
In the case $|\rho(0)=\ket{+}\bra{+}$, we find
@@
Finally, in the case $|\rho(0)=\ket{-}\bra{-}$, we find
@@
[…]
@@ Appendix C Analytical derivation of Eve accuracy for a projective measurement
The result above allows us to easily determine Eve accuracy in the case she performs a projective measurement at time $t^{*}$. Indeed, when the photon reaches Eve at time $t^{*}$, she also chooses randomly one of the two available bases. The probabilities above correspond to someone measuring in the right basis. However, if Eve does not, her probability of detecting the right state is still 1/2 as she can get each result with equal probability. Thus, the probability that Eve deduces the right state (i.e., the accuracy $A$) can be obtained from Eq. (43), which yields
@@ Appendix D Analytical derivation of Bob accuracy for a projective measurement (intercept-and-resend attack)
[…]
@@
[…]
@@
[…]
@@
Appendix E Information gain computation
Information gain from the homodyne currents